OpenStudy (anonymous):

HELP! Are these the equations: (Algebra) y= 2x + 31 y = 1x + 36 For their first birthday, Jaedan and Keegan's parents decided to plant trees that matched their sons' heights. Jaeden's tree was 31 in. tall, when it was planted and grows at the rate of 2 in. every yr. Keegan's tree was 36 in. when it was planted and it grows at the rate of 1 in. every yr. When will the two trees be at the same height?

OpenStudy (abb0t):

I don't like tree's. They creep me out.

OpenStudy (anonymous):

let after t yrs the tree will be same height then as per question 31+2t=36+t solving it 2t-t=36-31 or t=5 thus after 5 yrs the two trees be at the same height.